# Shotluck Holmes: A Family of Efficient Small-Scale Large Language Vision Models For Video Captioning and Summarization

### v1

@@ Abstract
Figure 1: A qualitative demo of the 3.1B model of Shotluck Holmes. Here, we demonstrate that our model is able to summarize data from many different parts of a long video using both visual and audio data.
1 Introduction
Over the last five years, the capability for machine learning models to intake, understand, and critically reason with natural language and visual data has exploded, primarily due to advances in model architecture Vaswani et al. (2017), compute capability, and a huge increase in available data. In particular, multi-modal large language models (LLMs), powered by the revolutionary transformer architecture Vaswani et al. (2017) have been able to achieve record-breaking understanding and reasoning capabilities on natural language and audiovisual understanding. Due to their resounding success in these fields OpenAI (2024), LLMs have also been at the forefront in the effort for AI models to understand video. Video is incredibly complex, since it combines dynamic movement in a visual medium with aural narration, sounds, text. These four aspects tend to constructively and destructively interfere with each other over the course of a given video. As such, it remains challenging for state-of-the-art (SOTA) language models to effectively comprehend and reason off of them. The current SOTA approach, Shot2Story20K Han et al. (2023), proposes a landmark new benchmark dataset that combines visual and auditory signals through a three-stage model pipeline. Then, they use this custom benchmark dataset to train a custom model architecture. This approach greatly improved performance on single-shot narration captioning, multi-shot video summarization, video Q&\&&A, and video retrieval, showing that embedding multiple sources of information is essential for language models to gain an improved understanding of these multi-modal inputs.
@@ 1 Introduction
2 Relevant Work
Large Multimodal Models.    Due to their impressive capabilities, large language models (LLMs) have garnered significant research interest in recent years Chowdhery et al. (2023); Chiang et al. ; Gunasekar et al. (2023); Zhang et al. (2024); Brown et al. (2020). This, combined with advancements in vision encoders Zhai et al. (2023); Sun et al. (2023), has led to some significant works in the multimodal Large Language Model field. Recent LLMs such as LLaVA Liu et al. (2024) and InstructBLIP Dai et al. (2024) leverage fine-tuning on existing LLM backbones with visual instruction tuning data to improve zero-shot performance and model alignment with human preferences. However, these models are rather large at 7B and 8.2B parameters respectively.
TinyLLaVA.    The TinyLLaVA framework provides analysis on exploiting various small-scale LLMs for LLVMs, utilizing the same fine-tuning approach on visual instruction tuning data. Their research Zhou et al. (2024) finds that SigLIP Zhai et al. (2023) yields better performance than CLIP when combined with small-scale LLMs of varying parameter ranges, including TinyLlama (1.1B) Zhang et al. (2024) and Phi-2 (2.7B) Li et al. (2023b). It was found that small-scale LLVMs (3.1B) can achieve better overall performance against existing 7B models including LLavA-1.5 and Qwen-VL on various evaluation benchmarks such as TextVQA Singh et al. (2019), SQA-I Lu et al. (2022), GQA Hudson and Manning (2019), VQAv2 Goyal et al. (2017), MMB Liu et al. (2023), MME Yin et al. (2023), LLaVA-W Liu et al. (2024), POPE Li et al. (2023a), and M-Vet Yu et al. (2023). However, the capabilities of TinyLLaVA are limited to just image-text-to-text generation.
Shot2Story20K.    Shot2Story20K Han et al. (2023) leverages HDvila100M Xue et al. (2022) (which contains abundant automatic speech recognition (ASR) content) and performs strict quality checking to collect a rich dataset of size 20023. A comparison of Shot2Story20K with pre-existing datasets (see Table 1) shows that Shot2Story20K outperforms its predecessors in quality. Recently, the authors have even released a 134K version of the dataset, placing this benchmark shot-level dataset far beyond the capabilities of its competitors.
Table 1: A high-level comparison of Shot2Story20K with pre-existing video-based datasets Han et al. (2023). M and G stand for manual and generated, respectively.
Dataset Annot.
Multi-shot
Video
Multi-event
Descriptions
Audio
Captions
Detailed
Summary
Summary
Length
# Vids.
Avg.
Duration
MSRVTT M - 10K 15s
ActivityNet Caps M 52.4 20K 3min
VideoStorytelling M 162.6 105 12.5min
Ego4D M - 10K 23min
YouCook2 M 67.8 2K 6min
VAST G 32.4 27M 5$\sim$30s
Shot2Story20K M+G 201.8 20K 16.7s
As mentioned, Shot2Story20K Han et al. (2023) combines visual and auditory signals by creating a dataset through a three-stage model. (1) TransNetV2 Souček and Lokoč (2020) separates videos into shots. (2) MiniGPT-4 Zhu et al. (2023) performs video captioning of individual shots and the results are manually checked by human annotators, who also write narration captions of individual shots. (3) GPT-4 OpenAI (2024) combines all individual shot captions and ASR to perform full video summarization with human prompting.
3 Method
3.1 Data-preprocessing
@@
3.2 Video to Tensor
In order for the LLM to process the video data, the video is first converted into a tensor and then fed into a vision encoder such as SigLip Zhai et al. (2023). To convert the video into a tensor, we experimented with two sampling methods inspired by the approach in LAVIS Li et al. (2022): uniform sampling and head-tail sampling. Head-tail sampling forces the random sampling to sample an equal number of frames from the first half of the video and the second half of the video. Sampled frames are then concatenated and fed into a vision encoder.
3.3 Shotluck Holmes Backbones
The Shot2Story20K model architecture uses multiple different language models to assist in the video summarization or Q&\&&A tasks. By leveraging more efficient small-scale LLVMs like TinyLLaVA, we hope to achieve strong performance on single-shot video captioning and multi-shot video summarization with significantly reduced computational complexity by replacing the entire Shot2Story20K model architecture with a single LLVM model Zhou et al. (2024). Because TinyLLaVA models are already pre-trained with vision encoders Zhai et al. (2023) and small-scale language models, our goal is to utilize Shot2Story20K’s dataset and fine-tune the smaller LLM to extract similar performance as compared to Shot2Story’s multi-step model.
Figure 2: Shotluck Holmes model architecture Zhou et al. (2024)
@@ 3.3 Shotluck Holmes Backbones
Table 2: Size of baseline models
@@
4 Experiments
@@
4.1 Single-Shot Video Captioning
The goal of single-shot video captioning is to generate descriptions for individual video shots (i.e. sections of a full video). The model first samples $N=120$ frames from a video shot using one of two sampling methods as described earlier. We chose $N=120$ based on the largest number of frames that we could feasiblty fit into our training hardware. These frames are then concatenated and fed into the vision encoder to produce visual tokens. The tokens are concatenated to a predetermined text prompt based on the type of LLM being used (see Appendix A), which is followed by the ASR text as additional context clues. Finally, this tensor is fed through the small-scale LLM of the given size to generate the caption for the video shot. Shot2Story20K discovered that applying ASR text as additional context cues in the text prompt results in slightly improved BLEU and METEOR metrics, with slight degradation in ROUGE. As a result, we follow this method and apply ASR text to our pipeline as well. The results are quite successful. For example, given the challenging shot below in Figure 3, the model generates the sentence: "In the video, a man in a black suit and tie is standing in front of a large screen displaying a boxing match. He is speaking into a microphone in front him."
Figure 3: The sampled frame for single-shot captioning
4.2 Multi-Shot Video Summarization
Multi-shot video summarization involves providing a rational summary describing a progression of events across different shots taken from the same video. We use a similar approach to that of single-shot captioning when training for multi-shot video summarization - our dataset included both single-shot and multi-shot video examples and the model was simultaneously finetuned to do both. Despite this lack of focus in finetuning, the model performs feasibly well as it is able to generate coherent summaries of multi-shot videos.
For example, for a video which contains different shots of a tablet computer and narration describing its specifications and features, the model generated the following summary: "The video is about a tablet computer that has a lot of features. It has a 64 gigabyte hard drive with two gigabytes of memory, a fast Intel Atom z3740 5F quad core data processor, and a detachable magnetic hinge that allows it to be easily connected to a keyboard. The tablet also has a latch list, which is a feature that allows you to quickly and easily engage and disengage the tablet from its keyboard".
4.3 Evaluation
Table 3: Performance of best models on single-shot video captioning
@@ 4.3 Evaluation
Figure 4: Single-shot video captioning results
@@
Table 4: Performance of best models on multi-shot video summarization
@@
Figure 5: Multi-shot video summarization results
@@
4.4 Qualitative Evaluation
Besides systematically evaluating on public benchmarks, we further qualitatively examined how well Shotluck Holmes performed on the video captioning tasks. As illustrated in Figure 1, Shotluck Holmes 3.1B was able to summarize information chronologically from the video. It was quite accurate at transcribing the narration and did not miss any key pieces of information. In Figure 3, the model was able to infer the context of the single-shot video (a boxing match) without explicitly being told that info in a narration.
4.5 Ablation Study: Sampling Methods
To understand the true underlying benefits of the new Shotluck-Holmes pipeline, we performed an ablation study to study the effects of different frame sampling strategies on model performance. Specifically, we aimed to investigate the impact of sampling methods on the performance of the model and the generalizability of our approach.
Table 5: Sampling Method Ablation Study on Shotluck-Holmes 1.5B using the Single-Shot Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (1.5B) w/ Uniform Sampling 9.3 25.3 36.3 68.9
Shotluck-Holmes (1.5B) w/ Head-Tail Sampling 8.2 25.5 36.5 65.0
Table 6: Sampling Method Ablation Study on Shotluck-Holmes 3.1B using the Single-Shot Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (3.1B) w/ Uniform Sampling 8.7 25.7 36.2 63.2
Shotluck-Holmes (3.1B) w/ Head-Tail Sampling 8.8 25.9 37.1 67.6
Table 7: Sampling Method Ablation Study on Shotluck-Holmes 1.5B using the Full Video Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (1.5B) w/ Uniform Sampling 6.48 21.3 40.2 144.3
Shotluck-Holmes (1.5B) w/ Head-Tail Sampling 6.40 21.4 41.0 151.9
Table 8: Sampling Method Ablation Study on Shotluck-Holmes 3.1B using the Full Video Task
Model BLEU METEOR ROUGE CIDER
Shotluck-Holmes (3.1B) w/ Uniform Sampling 7.67 23.2 43 152.3
Shotluck-Holmes (3.1B) w/ Head-Tail Sampling 7.60 23.0 42.4 157.9
To investigate this, we studied two common video frame sampling techniques. The first technique, known as uniform sampling, gives each frame a equal probability to be selected. The dataloader then selected $N$ frames at random (in our experiments, we kept $N=120$, for consistency). In comparison, head-tail sampling disproportionately weights the likelihood of drawing frames from the beginning and end of a given shot or video to get a total of $N$ sampled frames. The results of each, evaluated using the standard four metrics, are in the four tables above.
As demonstrated by the studies, there is no significant trend observable between the two sampling techniques. In some models and tasks, head-tail sampling performed better, while in others, uniform sampling performed better. Nevertheless, the two sampling methods resulted not only in comparable results, but also both soundly beat the SOTA Shot2Story20K model, demonstrating that our approach is generalizable across sampling methods.
5 Conclusion
In this paper, we propose Shotluck-Holmes, a family of efficient models that achieve state-of-the-art performance on shot-level and full-length video understanding. This result is achieved by combining Shot2Story’s Han et al. (2023) multi-model pipeline, which integrates shot-level video annotations with audio-visual elements, with small-scale LLMs like TinyLLaVA. Furthermore, we demonstrate that these smaller models are able to achieve generalization capabilities on video captioning and summarization tasks that are competitive with larger models, suggesting that such a task is executable even on limited hardware or on edge devices.
@@ 5 Conclusion
6 Acknowledgements
We would like to thank Dr. Wei Xu and her teaching assistants (Anton Lavrouk, Marcus Ma, Andrew Li) for the course CS 7650: Natural Langauge Processing at Georgia Tech, from which this project was inspired. In addition, we would like to extend a major thank you to PACE @ Georgia Tech PACE (2017) for providing the plentitude of computational resources and GPUs needed to train our models.
@@ Appendix B Hyperparameters
Table 9: Hyperparameters for Shotluck Holmes training
@@
Table 10: Evaluation Parameter Settings

### v2

@@ Abstract.
Deep Learning, Multimodal Models, Large Language Models, Machine Learning, Natural Language Processing, Vision, Vision-Language Models
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2nd International Workshop on Deep Multimodal Generation and Retrieval; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 2nd International Workshop on Deep Multimodal Generation and Retrieval (MMGR ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3689091.3690086isbn: 979-8-4007-1202-9/24/10conference: Proceedings of the 2nd International Workshop on Deep Multimodal Generation and Retrieval; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 2nd International Workshop on Deep Multimodal Generation and Retrieval (MMGR ’24), October 28-November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-1202-9/24/10submissionid: mmgr1904ccs: Computing methodologies Neural networkscopyright: none
1. Introduction
Over the last five years, the capability for machine learning models to intake, understand, and critically reason with language and visual data has exploded, primarily due to advances in model architecture (Vaswani et al., 2017), compute capability, and a huge increase in available data. In particular, multi-modal large language models (LLMs), powered by the revolutionary transformer architecture (Vaswani et al., 2017) have been able to achieve record-breaking understanding and reasoning capabilities on natural language and audiovisual understanding. Due to their resounding success in these fields (OpenAI, 2024), LLMs have also been at the forefront in building intelligent agents to understand video. Video is incredibly complex, since it combines dynamic movement in a visual medium with aural narration, sounds, text. These four aspects tend to constructively and destructively interfere with each other over the course of a given video. As such, it remains challenging for state-of-the-art (SOTA) language models to effectively comprehend and reason off of them. The current SOTA approach, Shot2Story20K (Han et al., 2023), proposes a landmark new benchmark dataset that combines visual and auditory signals through a three-stage model pipeline. Then, they use this custom benchmark dataset to train a custom model architecture. This approach greatly improved performance on single-shot narration captioning, multi-shot video summarization, video Q&\&&A, and video retrieval, showing that embedding multiple sources of information is essential for language models to gain an improved understanding of these multi-modal inputs.
@@ 1. Introduction
2. Relevant Work
Large Multimodal Models.    Due to their impressive capabilities, large language models (LLMs) have garnered significant research interest in recent years (Chowdhery et al., 2023; Chiang et al., [n. d.]; Gunasekar et al., 2023; Zhang et al., 2024; Brown et al., 2020). This, combined with advancements in vision encoders (Zhai et al., 2023; Sun et al., 2023), has led to some significant works in the multimodal Large Language Model field. Recent LLMs such as LLaVA (Liu et al., 2024) and InstructBLIP (Dai et al., 2024) leverage fine-tuning on existing LLM backbones with visual instruction tuning data to improve zero-shot performance and model alignment with human preferences. However, these models are rather large at 7B and 8.2B parameters respectively.
TinyLLaVA.    The TinyLLaVA framework provides analysis on exploiting various small-scale LLMs for LLVMs, utilizing the same fine-tuning approach on visual instruction tuning data. Their research (Zhou et al., 2024) finds that SigLIP (Zhai et al., 2023) yields better performance than CLIP when combined with small-scale LLMs of varying parameter ranges, including TinyLlama (1.1B) (Zhang et al., 2024) and Phi-2 (2.7B) (Li et al., 2023a). It was found that small-scale LLVMs (3.1B) can achieve better overall performance against existing 7B models including LLavA-1.5 and Qwen-VL on various evaluation benchmarks such as TextVQA (Singh et al., 2019), SQA-I (Lu et al., 2022), GQA (Hudson and Manning, 2019), VQAv2 (Goyal et al., 2017), MMB (Liu et al., 2023), MME (Yin et al., 2023), LLaVA-W (Liu et al., 2024), POPE (Li et al., 2023b), and M-Vet (Yu et al., 2023). However, the capabilities of TinyLLaVA are limited to just image-text-to-text generation.
Shot2Story20K.    Shot2Story20K (Han et al., 2023) leverages HDvila100M (Xue et al., 2022) (which contains abundant automatic speech recognition (ASR) content) and performs strict quality checking to collect a rich dataset of size 20023. Recently, the authors have even released a 134K version of the dataset, placing this benchmark shot-level dataset far beyond the capabilities of its competitors.
As mentioned, Shot2Story20K (Han et al., 2023) combines visual and auditory signals by creating a dataset through a three-stage model. (1) TransNetV2 (Souček and Lokoč, 2020) separates videos into shots. (2) MiniGPT-4 (Zhu et al., 2023) performs video captioning of individual shots and the results are manually checked by human annotators, who also write narration captions of individual shots. (3) GPT-4 (OpenAI, 2024) combines all individual shot captions and ASR to perform full video summarization with human prompting.
3. Method
3.1. Data-Preprocessing
@@
3.2. Video to Tensor Conversion
In order for the LLM to process the video data, the video is first converted into a tensor and then fed into a vision encoder such as SigLip (Zhai et al., 2023). To convert the video into a tensor, we experimented with two sampling methods inspired by the approach in LAVIS (Li et al., 2022): uniform sampling and head-tail sampling. Head-tail sampling forces the random sampling to sample an equal number of frames from the first half of the video and the second half of the video. Sampled frames are then concatenated and fed into a vision encoder.
3.3. Shotluck Holmes Backbone
The Shot2Story20K model architecture uses multiple different language models to assist in the video summarization or Q&\&&A tasks. By leveraging more efficient small-scale LLVMs like TinyLLaVA, we hope to achieve strong performance on single-shot video captioning and multi-shot video summarization with significantly reduced computational complexity by replacing the entire Shot2Story20K model architecture with a single LLVM model (Zhou et al., 2024). Because TinyLLaVA models are already pre-trained with vision encoders (Zhai et al., 2023) and small-scale language models, our goal is to utilize Shot2Story20K’s dataset and fine-tune the smaller LLM to extract similar performance as compared to Shot2Story’s multi-step model.
Figure 1. Shotluck Holmes model architecture (Zhou et al., 2024)
@@ 3.3. Shotluck Holmes Backbone
Table 1. Size of baseline models
@@
3.4. Single-Shot Video Captioning
The goal of single-shot video captioning is to generate descriptions for individual video shots (i.e. sections of a full video). The model first samples $N=120$ frames from a video shot using one of two sampling methods as described earlier. We chose $N=120$ based on the largest number of frames that we could feasiblty fit into our training hardware. These frames are then concatenated and fed into the vision encoder to produce visual tokens. The tokens are fed into a MLP and then concatenated to a predetermined text prompt based on the type of LLM being used (see Appendix A) with ASR text as additional context clues. Finally, this tensor is fed through the small-scale LLM of the given size to generate the caption for the video shot.
3.5. Multi-shot Video Summarization
Multi-shot video summarization involves providing a rational summary describing a progression of events across different shots taken from the same video. We follow the same approach as the single-shot video captioning: we sample the same number of frames, except this time we sample from the entire video. ASR information is also retrieved from the entire video.
4. Experiments
@@
4.1. Single-Shot Video Captioning
With the addition of context clues and ASR, the results are quite successful. For example, given the challenging shot below in Figure 2, the model generates the sentence: ”In the video, a man in a black suit and tie is standing in front of a large screen displaying a boxing match. He is speaking into a microphone in front him.”
Figure 2. The sampled frame for single-shot captioning
4.2. Multi-Shot Video Summarization
Our dataset included both single-shot and multi-shot video examples and the model was simultaneously finetuned to do both. Despite this lack of focus in finetuning, the model performs feasibly well as it is able to generate coherent summaries of multi-shot videos.
For example, for a video which contains different shots of a tablet computer and narration describing its specifications and features, the model generated the following summary: ”The video is about a tablet computer that has a lot of features. It has a 64 gigabyte hard drive with two gigabytes of memory, a fast Intel Atom z3740 5F quad core data processor, and a detachable magnetic hinge that allows it to be easily connected to a keyboard. The tablet also has a latch list, which is a feature that allows you to quickly and easily engage and disengage the tablet from its keyboard”.
4.3. Evaluation
Table 2. Performance of best models on single-shot video captioning
@@ 4.3. Evaluation
Figure 3. Single-shot video captioning results
@@
Table 3. Performance of best models on multi-shot video summarization
@@
Figure 4. Multi-shot video summarization results
@@
4.4. Qualitative Evaluation
Besides systematically evaluating on public benchmarks, we further qualitatively examined how well Shotluck Holmes performed on the video captioning tasks. Shotluck Holmes 3.1B was able to summarize information chronologically from the video. It was quite accurate at transcribing the narration and did not miss any key pieces of information. In Figure 2, the model was able to infer the context of the single-shot video (a boxing match) without explicitly being told that info in a narration.
5. Conclusion
In this paper, we propose Shotluck-Holmes, a family of efficient models that achieve state-of-the-art performance on shot-level and full-length video understanding. This result is achieved by combining Shot2Story’s (Han et al., 2023) multi-model pipeline, which integrates shot-level video annotations with audio-visual elements, with small-scale LLMs like TinyLLaVA. Furthermore, we demonstrate that these smaller models are able to achieve generalization capabilities on video captioning and summarization tasks that are competitive with larger models, suggesting that such a task is executable even on limited hardware or on edge devices.
@@ Appendix B Hyperparameters
Table 4. Hyperparameters for Shotluck Holmes training
@@
Table 5. Evaluation Parameter Settings